OpenStudy (anonymous):

Write the arithmetic sequence that has 3 arithmetic means between the following two fractions

OpenStudy (anonymous):

\[-\frac{1}{4} and -\frac{19}{12}\]

OpenStudy (binary3i):

arithmetic mean between x and y is (x+y)/2 = b AM between x and b is (x+b)/2=a AM between b and y is (b+y)/2=c now the common difference of the sequence is a- x Tth term= X + (a-x)(t -1)

OpenStudy (anonymous):

I'm still pretty confused but \[a=-\frac{7}{12}\] \[b=-\frac{11}{12}\] \[c=-\frac{5}{4}\] Are these all right? I don't know what to do now

OpenStudy (binary3i):

\[x=-1/4 , y= -19/12\]

OpenStudy (anonymous):

That's what I had as x and y. Are my answers wrong?